A skydiver with a mass of 68.96 kg dives out of a perfectly good airplane and plummets toward the Earth below. After falling for several seconds, the diver's acceleration is measured at 4.05 m/s2. What is the magnitude of the force of air resistance at this moment?
1
Expert's answer
2017-01-26T13:32:15-0500
ma=mg-F_res,
F_res=m (g-a)=68.96 (9.8-4.05)=396.52 N